# Spatial Sign based Principal Component Analysis for High Dimensional Data

### v1

@@ 3.1 Combinatoric Program
In this section, we consider sparse settings: $\lambda_{1}({\bf\Sigma})$ is distinct and $||\boldsymbol{u}_{1}({\bf\Sigma})||_{0}\leq s<d\wedge n$. For any matrix $\mathbf{M}\in\mathbb{R}^{d\times d}$, we define the best $s$-sparse vector approximating $\boldsymbol{u}_{1}(\mathbf{M})$ as
@@ 3.2 Computationally Efficient program
We adopt the truncated power algorithm proposed by Yuan and Zhang (2013) to solve the optimization problem (7). For any vector $\boldsymbol{v}\in\mathbb{R}^{d}$ and an index set $J\subset\{1,\ldots,d\}$, we define the truncation function $\operatorname{TRC}(\cdot,\cdot)$ to be
$\operatorname{TRC}(\boldsymbol{v},J):=\left(v_{1}\cdot\mathbb{I}(1\in J),\ldots,v_{d}\cdot\mathbb{I}(d\in J)\right)^{T}$
where $\mathbb{I}(\cdot)$ is the indicator function. Algorithm 1 shows the detail procedures of our proposed SSPCA procedure.
@@ 4.1 Estimating Leading Eigenvector
We first consider estimating the leading eigenvector of the covariance matrix ${\bf\Sigma}$. We consider the similar model for $\bf\Sigma$ as citehan2018eca, i.e.
@@
    Multivariate normal distribution. $\boldsymbol{X}\sim N(\boldsymbol{\theta},{\bf\Sigma})$.
@@
To evaluate the dependence of the estimation accuracy of the SSPCA estimator on the triplet $(n,d,s)$, we conducted experiments with varying values of $d$, $s$, and sample size $n$. Specifically, we considered $d=100,200,300$, $s_{1}=5,10,20$, and varying sample sizes $n$. The results, presented in Figure 2, show the curves of averaged distances between the estimates and true parameters. In these experiments, we set the number of selected features equal to the true parameter $s$. Our findings indicate that the averaged distance between $\boldsymbol{v}_{1}$ and $\hat{\boldsymbol{v}}_{1}$ approaches zero as the sample size increases, which demonstrates the consistency of our proposed SSPCA methods. This consistency is an important characteristic of any estimator, as it indicates that the estimates produced by the method will be increasingly accurate as more data is available. Additionally, we observed that all the curves in Figure 2 almost overlap with each other when the average distances are plotted against $\log d/n$. This observation is consistent with the results presented in Theorem ??, which suggests that the effective sample size is $n/\log d$ when controlling the prediction accuracy of the eigenvectors. This finding highlights the importance of considering the relationship between $n$ and $d$ when evaluating the performance of SSPCA. Specifically, it suggests that as the dimension $d$ increases, the sample size $n$ must also increase in order to maintain a given level of prediction accuracy.
@@ 5 Real Data Analysis
In this section, we apply three methodologies—Thresholding Pursuit (TP), Exponential Component Analysis (ECA), and Sparse and Structured Principal Component Analysis (SSPCA)—to analyze the Standard & Poor’s 500 (S&P 500) index. To account for the dynamic nature of the index’s composition over time, we compiled monthly returns for all securities included in the S&P 500 from January 2005 to November 2018 ($n=165$). Given the evolving nature of the index, we focused on a consistent subset of $d=374$ securities that were present throughout this entire period. As demonstrated in Liu et al. (2023), stock returns exhibit non-Gaussian, heavy-tailed characteristics, which necessitate the use of robust statistical procedures. For simplicity, our analysis considers only the first two principal components.
@@ 6 Conclusion
7 Appendix
7.1 Appendix A: Fantope Projection
@@
Next, we conduct simulation studies to compare the two distinct initial estimators. SSPCA denotes the proposed method utilizing the eigenvector as the initial estimator, whereas SSPCA-FP refers to the proposed method incorporating Fantope Projection. In this analysis, we focus solely on the estimation of the first eigenvector. Figure 9 presents the outcomes of these two approaches. Our observations reveal that when the number of selected features matches the true number, SSPCA-FP performs similar to SSPCA. However, when the number of selected features exceeds the true parameters, SSPCA-FP exhibits smaller averaged distances compared to SSPCA. We also compare these two methods with the averaged distances between the estimates and true parameters of the leading eigenvector with estimated number of selected features $\hat{s}$. Table 3 reveals the simulation results with the same settings as Table 1. Overall, the performance of these two initial estimators is quite comparable. SSPCA-FP exhibits slightly smaller averaged distances when compared to the estimator $\hat{s}$. Therefore, if computational time is not a constraint, we recommend using SSPCA-FP as the primary choice.
Figure 9: Curves of averaged distances between the estimates and true parameters with different number of selected features on estimating leading eigenvectors.
@@
7.2 Appendix B: Proof of Theorems
7.2.1 Some useful lemmas
The accuracies of constant and linear approximations of function $|\mathbf{y}-\mu|^{-1}(\mathbf{y}-\mu)$ of $\mu$ are given by 1 .
LEMMA 7.1
@@ 7.2.1 Some useful lemmas
LEMMA 7.2
The accuracies of constant, linear and quadratic approximations of the function $\boldsymbol{\mu}\mapsto\|\mathbf{z}-\boldsymbol{\mu}\|$ can be given by
@@
    $|\|\mathbf{z}-\boldsymbol{\mu}\|-\|\mathbf{z}\||\leq\|\boldsymbol{\mu}\|$,
@@
    $\left|\|\mathbf{z}-\boldsymbol{\mu}\|-\|\mathbf{z}\|+\mathbf{u}^{T}\boldsymbol{\mu}\right|\leq 2r^{-1}\|\boldsymbol{\mu}\|^{2}$,
@@
    $\left|\|\mathbf{z}-\boldsymbol{\mu}\|-\|\mathbf{z}\|+\mathbf{u}^{T}\boldsymbol{\mu}-\boldsymbol{\mu}^{T}(2r)^{-1}\left[\mathbf{I}_{p}-\mathbf{u}\mathbf{u}^{T}\right]\boldsymbol{\mu}\right|\leq cr^{-1-\delta}\|\boldsymbol{\mu}\|^{2+\delta}$ for all $10<\delta<$ 1,
where $\mathbf{z}=r\mathbf{u},r=\|\mathbf{z}\|,\mathbf{u}=\|\mathbf{z}\|^{-1}\mathbf{z}$ and the constant $c$ does not depend on $\mathbf{z}$ or $\boldsymbol{\mu}$.
@@
$L(\boldsymbol{\theta})=\sum_{i=1}^{n}\|\boldsymbol{X}_{i}-\boldsymbol{\mu}-\boldsymbol{\theta}\|-\sum_{i=1}^{n}\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|$
and $\hat{\boldsymbol{\theta}}=\arg\max L(\boldsymbol{\theta})$. Then $\hat{\boldsymbol{\theta}}=\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}$. Next, we proof that for any $\epsilon>0$, there exists $C>0$ such that
@@
$P\left(\|\hat{\boldsymbol{\theta}}\|\leq C\zeta_{1}^{-1}n^{-1/2}\right)\geq 1-\epsilon,$
which means $\hat{\boldsymbol{\theta}}=O_{p}(\zeta_{1}^{-1}n^{-1/2})$. By Lemma 7.2, we have
@@
$\leq\frac{4{\rm Var}(\zeta_{1}^{-1}r_{i}^{-1}[1-(\boldsymbol{U}_{i}^{T}\boldsymbol{u})^{2}])}{n(1-\boldsymbol{u}^{T}{\bf S}\boldsymbol{u})}\leq\frac{4\zeta}{n\psi}\leq\frac{\epsilon}{3}$
@@
by condition (A1) for sufficient large $n$. So $A_{3}\geq\frac{cC^{2+\delta}}{2\zeta_{1}}E(\nu_{i}^{1+\delta})n^{-\frac{1}{2}\delta}$. Thus, at least probability $1-\epsilon$, we have
@@
$\sum_{i=1}^{n}U(\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}})=\sum_{i=1}^{n}\frac{\boldsymbol{X}_{i}-\boldsymbol{\mu}-\hat{\boldsymbol{\theta}}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}-\hat{\boldsymbol{\theta}}\|}=\sum_{i=1}^{n}\frac{\boldsymbol{U}_{i}-r_{i}^{-1}\hat{\boldsymbol{\theta}}}{(1+2r_{i}^{-1}\boldsymbol{U}_{i}^{T}\hat{\boldsymbol{\theta}}+r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2})^{1/2}}$
@@
$\left\{n^{-1}\sum_{i=1}^{n}r_{i}^{-1}+O_{p}\left(\zeta_{1}n^{-1/2}\right)\right\}\hat{\boldsymbol{\theta}}-\left\{n^{-1}\sum_{i=1}^{n}r_{i}^{-1}\mathbf{U}_{i}\mathbf{U}_{i}^{T}\right\}\hat{\boldsymbol{\theta}}=\left(n^{-1}\sum_{i=1}^{n}\mathbf{U}_{i}\right)\left(1+O_{p}\left(n^{-1}\right)\right).$
Then, we obtained that
$\hat{\boldsymbol{\theta}}=\left(\zeta_{1}+O_{p}\left(\zeta_{1}n^{-1/2}\right)\right)^{-1}\left(n^{-1}\sum_{i=1}^{n}\mathbf{U}_{i}\right)\left\{1+O_{p}\left(n^{-1}\right)\right\}=n^{-1}\zeta_{1}^{-1}\sum_{i=1}^{n}\mathbf{U}_{i}+O_{p}\left(\zeta_{1}^{-1}n^{-1}\right)$
7.2.2 Proof of Theorem 2.1
@@
Thus, according to Lemma 7.3, we have
@@
By lemma 7.4, we have, there exist a positive constant $C_{\alpha}$ such that $\left\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\right\|_{2}^{2}\leq C_{\alpha}^{2}\zeta_{1}^{-2}n^{-1}$ with probability larger than $61-\alpha/$ 6. Additionally, by Chebyshev inequality,
@@
By Lemma 7.4, we can write $\hat{\boldsymbol{\theta}}=\zeta_{1}^{-1}\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{U}_{i}+\zeta_{1}^{-1}\varrho$, where $\|\varrho\|_{2}=O_{p}(n^{-1})$. So
@@
So $P\left(\|B_{1}\|\geq 3\sqrt{2}\alpha^{-1/2}n^{-1}\sigma_{v}\|{\bf S}\|_{2}\right)\leq\frac{\alpha}{18}.$ Additionally,
@@
7.2.4 Proof of Theorem 3.1
LEMMA 7.5
@@
Proof of Theorem 3.1: Note that $U(\boldsymbol{X})$ has the same distribution as $S(X)=\frac{\boldsymbol{X}-\tilde{\boldsymbol{X}}}{\|\boldsymbol{X}-\tilde{\boldsymbol{X}}\|}$ where $\boldsymbol{X},\tilde{\boldsymbol{X}}\sim EC_{d}(\boldsymbol{\mu},{\bf\Sigma},\xi)$ and are independent. By Lemma B.4 in Han and Liu (2018), for any $\boldsymbol{v}=\left(v_{1},\ldots,v_{d}\right)^{T}\in\mathbb{S}^{d-1}$, Equation (10) holds with
@@
where $\left(Y_{1},\ldots,Y_{d}\right)^{T}\sim N_{d}\left(\mathbf{0},\mathbf{I}_{d}\right)$ is standard Gaussian. Thus, by Lemma 7.5, when $(s\log(ed/s)+\log(1/\alpha))/n\rightarrow 0$, with probability at least $1-2\alpha$, we have
$\|\widetilde{\mathbf{S}}-\mathbf{S}\|_{2,s}\leq$ $C_{0}\left(\sup_{\boldsymbol{v}\in\mathbb{S}^{d-1}}2\left\|\boldsymbol{v}^{T}U(\boldsymbol{X})\right\|_{\psi_{2}}^{2}+\|\mathbf{S}\|_{2}\right)\sqrt{\frac{s(3+\log(d/s))+\log(1/\alpha)}{n}}.$
Taking the same procedure as Lemma 7.4, we can have $\|\hat{\boldsymbol{\theta}}\|_{2,s}=O_{p}(\sqrt{\frac{s}{d}}\zeta_{1}^{-1}n^{-1/2})$. So taking the same procedure as the proof of Theorem 2.1, we have, there exist some positive constant $C_{1}>0$,
@@
for sufficient large $n$ and with probability larger than $1-\alpha$. Then, by the triangle inequality, we obtain the first result. Specially, if $\operatorname{rank}(\boldsymbol{\Sigma})=q$ and $\left\|\boldsymbol{u}_{1}(\boldsymbol{\Sigma})\right\|_{0}\leq s$, by Theorem 4.2 in citehan2018eca, we have

### v2

@@ 3.1 Combinatoric Program
In this section, we consider sparse settings: $\lambda_{1}({\bf\Sigma})$ is distinct and $\|\boldsymbol{u}_{1}({\bf\Sigma})\|_{0}\leq s<d\wedge n$. For any matrix $\mathbf{M}\in\mathbb{R}^{d\times d}$, we define the best $s$-sparse vector approximating $\boldsymbol{u}_{1}(\mathbf{M})$ as
@@ 3.2 Computationally Efficient program
We adopt the truncated power algorithm proposed by Yuan and Zhang (2013) to solve the optimization problem (7). For any vector $\boldsymbol{v}\in\mathbb{R}^{d}$ and an index set $J\subset\{1,\ldots,d\}$, we define the truncation function $\operatorname{TRC}(\cdot,\cdot)$ to be $\operatorname{TRC}(\boldsymbol{v},J):=\left(v_{1}\cdot\mathbb{I}(1\in J),\ldots,v_{d}\cdot\mathbb{I}(d\in J)\right)^{T}$ where $\mathbb{I}(\cdot)$ is the indicator function. Algorithm 1 shows the detail procedures of our proposed SSPCA procedure.
@@ 4.1 Estimating Leading Eigenvector
We first consider estimating the leading eigenvector of the covariance matrix ${\bf\Sigma}$. We consider the similar model for $\bf\Sigma$ as Han and Liu (2018), i.e.
@@
    Multivariate normal distribution. $\boldsymbol{X}\sim N({\boldsymbol{0}},{\bf\Sigma})$.
@@
To evaluate the dependence of the estimation accuracy of the SSPCA estimator on the triplet $(n,d,s)$, we conducted experiments with varying values of $d$, $s$, and sample size $n$. Specifically, we considered $d=100,200,300$, $s_{1}=5,10,20$, and varying sample sizes $n$. The results, presented in Figure 2, show the curves of averaged distances between the estimates and true parameters. In these experiments, we set the number of selected features equal to the true parameter $s$. Our findings indicate that the averaged distance between $\boldsymbol{v}_{1}$ and $\hat{\boldsymbol{v}}_{1}$ approaches zero as the sample size increases, which demonstrates the consistency of our proposed SSPCA methods. This consistency is an important characteristic of any estimator, as it indicates that the estimates produced by the method will be increasingly accurate as more data is available. Additionally, we observed that all the curves in Figure 2 almost overlap with each other when the average distances are plotted against $\log d/n$. This observation is consistent with the results presented in Corollary 2.1, which suggests that the effective sample size is $n/\log d$ when controlling the prediction accuracy of the eigenvectors. This finding highlights the importance of considering the relationship between $n$ and $d$ when evaluating the performance of SSPCA. Specifically, it suggests that as the dimension $d$ increases, the sample size $n$ must also increase in order to maintain a given level of prediction accuracy.
@@ 5 Real Data Analysis
5.1 S&P 500 Index Stock Data
In this subsection, we apply three methodologies—Thresholding Pursuit (TP), Exponential Component Analysis (ECA), and Sparse and Structured Principal Component Analysis (SSPCA)—to analyze the Standard & Poor’s 500 (S&P 500) index. To account for the dynamic nature of the index’s composition over time, we compiled monthly returns for all securities included in the S&P 500 from January 2005 to November 2018 ($n=165$). Given the evolving nature of the index, we focused on a consistent subset of $d=374$ securities that were present throughout this entire period. As demonstrated in Liu et al. (2023), stock returns exhibit non-Gaussian, heavy-tailed characteristics, which necessitate the use of robust statistical procedures. For simplicity, our analysis considers only the first two principal components.
@@ 5.1 S&P 500 Index Stock Data
5.2 MNIST Dataset
In this subsection, we apply three methodologies to analyze the MNIST dataset (LeCun et al., 2002). The MNIST dataset comprises 60,000 grayscale images of handwritten digits ranging from zero to nine. Each image is $28\times 28$ pixels in size and is labeled with its corresponding digit. For our analysis, we construct the training matrix using the first 660 samples of the digit “1” and the first 33 samples of the digit “7,” resulting in a $693\times 784$ matrix. The remaining samples labeled as “1” and “7” constitute the test set, forming a $12,314\times 784$ matrix. The training data are standardized (zero mean, unit variance), and the same parameters are used to standardize the test data.
For simplicity, we focus on the first two principal components in each method. Using a parameter selection procedure over the candidate set $k\in\{300,350,400,450,500,550,600\}$, we determine the optimal values for the first and second principal components for each of the three methods. Subsequently, the two principal components derived from each method are used to train a Support Vector Machine (SVM) classifier on the training set. The trained model is then evaluated on the test set, yielding the following classification accuracies: 0.4939 for TP, 0.8221 for ECA, and 0.8589 for SSPCA. Figure 9 presents the scatter plots of the first (PC1) and second (PC2) principal components obtained from TP, ECA, and SSPCA, respectively. In each plot, red dots indicate potential outliers with strong leverage effects. A threshold of 0.02 is employed to identify such influential data points. Our analysis revealed that 14 data points have strong leverage influence for the TP method, 3 for the ECA method, and 0 for the SSPCA method. These results demonstrate the robustness of SSPCA in mitigating the influence of outliers.
Additionally, we conduct 100 simulation experiments. In each experiment, 660 samples labeled “1” and 33 samples labeled “7” are randomly drawn from the training set to form the training data, with the remaining “1” and “7” samples used for testing. The same standardization and parameter selection procedure as described above is applied in each experiment. We compute the average classification accuracy across the 100 runs, yielding the following results: 0.4939 for TP, 0.7154 for ECA, and 0.8074 for SSPCA.
These results demonstrate the advantages of the proposed SSPCA method. It consistently outperforms TP and ECA in classification accuracy and shows greater robustness to outliers, as evidenced by the absence of high-leverage points. This indicates that SSPCA provides more stable and reliable representations in high-dimensional settings.
@@ 6 Conclusion
Supplemental Material of “Spatial Sign based Principal Component Analysis for High Dimensional Data”
S1 Appendix A: Fantope Projection
@@
Next, we conduct simulation studies to compare the two distinct initial estimators. SSPCA denotes the proposed method utilizing the eigenvector as the initial estimator, whereas SSPCA-FP refers to the proposed method incorporating Fantope Projection. In this analysis, we focus solely on the estimation of the first eigenvector. Figure S10 presents the outcomes of these two approaches. Our observations reveal that when the number of selected features matches the true number, SSPCA-FP performs similar to SSPCA. However, when the number of selected features exceeds the true parameters, SSPCA-FP exhibits smaller averaged distances compared to SSPCA. We also compare these two methods with the averaged distances between the estimates and true parameters of the leading eigenvector with estimated number of selected features $\hat{s}$. Table S3 reveals the simulation results with the same settings as Table 1. Overall, the performance of these two initial estimators is quite comparable. SSPCA-FP exhibits slightly smaller averaged distances when compared to the estimator $\hat{s}$. Therefore, if computational time is not a constraint, we recommend using SSPCA-FP as the primary choice.
Figure S10: Curves of averaged distances between the estimates and true parameters with different number of selected features on estimating leading eigenvectors.
@@
Table S3: The averaged distances between the estimates and true parameters of the leading eigenvector with estimated number of selected features $\hat{s}$ and the true number of selected features $s$.
@@
S2 Appendix B: Proof of Theorems
S2.1 Some useful lemmas
The accuracies of constant and linear approximations of function $|\mathbf{y}-\mu|^{-1}(\mathbf{y}-\mu)$ of $\mu$ are given by Oja (2010).
LEMMA S2.1
@@
LEMMA S2.2
The accuracies of constant, linear and quadratic approximations of the function $\boldsymbol{\mu}\mapsto\|\mathbf{z}-\boldsymbol{\mu}\|{{}_{2}}$ can be given by
@@
    $|\|\mathbf{z}-\boldsymbol{\mu}\|{{}_{2}}-\|\mathbf{z}\|{{}_{2}}|\leq\|\boldsymbol{\mu}\|{{}_{2}}$,
@@
    $\left|\|\mathbf{z}-\boldsymbol{\mu}\|{{}_{2}}-\|\mathbf{z}\|{{}_{2}}+\mathbf{u}^{T}\boldsymbol{\mu}\right|\leq 2r^{-1}\|\boldsymbol{\mu}\|{{}_{2}}^{2}$,
@@
    $\left|\|\mathbf{z}-\boldsymbol{\mu}\|{{}_{2}}-\|\mathbf{z}\|{{}_{2}}+\mathbf{u}^{T}\boldsymbol{\mu}-\boldsymbol{\mu}^{T}(2r)^{-1}\left[\mathbf{I}_{p}-\mathbf{u}\mathbf{u}^{T}\right]\boldsymbol{\mu}\right|\leq cr^{-1-\delta}\|\boldsymbol{\mu}\|{{}_{2}}^{2+\delta}$ for all $10<\delta<$ 1,
where $\mathbf{z}=r\mathbf{u},r=\|\mathbf{z}\|{{}_{2}},\mathbf{u}=\|\mathbf{z}\|{{}_{2}}^{-1}\mathbf{z}$ and the constant $c$ does not depend on $\mathbf{z}$ or $\boldsymbol{\mu}$.
@@
$L(\boldsymbol{\theta})=\sum_{i=1}^{n}\|\boldsymbol{X}_{i}-\boldsymbol{\mu}-\boldsymbol{\theta}\|{{}_{2}}-\sum_{i=1}^{n}\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|{{}_{2}}$
and $\hat{\boldsymbol{\theta}}=\arg\min L(\boldsymbol{\theta})$. Then $\hat{\boldsymbol{\theta}}=\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}$. Next, we proof that for any $\epsilon>0$, there exists $C>0$ such that
@@
$P\left(\|\hat{\boldsymbol{\theta}}\|{{}_{2}}\leq C\zeta_{1}^{-1}n^{-1/2}\right)\geq 1-\epsilon,$
which means $\hat{\boldsymbol{\theta}}=O_{p}(\zeta_{1}^{-1}n^{-1/2})$. By Lemma S2.2, we have
@@
$\leq\frac{16{\rm Var}(\zeta_{1}^{-1}r_{i}^{-1}[1-(\boldsymbol{U}_{i}^{T}\boldsymbol{u})^{2}])}{n(1-\boldsymbol{u}^{T}{\bf S}\boldsymbol{u})^{2}}\leq\frac{16\zeta}{n\psi^{2}}\leq\frac{\epsilon}{3}$
@@
by condition (A1) for sufficient large $n$. So $A_{3}\geq-\frac{cC^{2+\delta}}{2\zeta_{1}}E(\nu_{i}^{1+\delta})n^{-\frac{1}{2}\delta}$. Thus, at least probability $1-\epsilon$, we have
@@
$\sum_{i=1}^{n}U(\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}})=\sum_{i=1}^{n}\frac{\boldsymbol{X}_{i}-\boldsymbol{\mu}-\hat{\boldsymbol{\theta}}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}-\hat{\boldsymbol{\theta}}\|{{}_{2}}}=\sum_{i=1}^{n}\frac{\boldsymbol{U}_{i}-r_{i}^{-1}\hat{\boldsymbol{\theta}}}{(1-2r_{i}^{-1}\boldsymbol{U}_{i}^{T}\hat{\boldsymbol{\theta}}+r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2})^{1/2}}$
@@
$n^{-1}\sum_{i=1}^{n}\left(\boldsymbol{U}_{i}-r_{i}^{-1}\hat{\boldsymbol{\theta}}\right)\left(1-2r_{i}^{-1}\boldsymbol{U}_{i}^{\top}\hat{\boldsymbol{\theta}}+r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2}\right)^{-1/2}=0.$
By the Taylor expansion, the above equation can be rewritten as
$n^{-1}\sum_{i=1}^{n}\left(\boldsymbol{U}_{i}-r_{i}^{-1}\hat{\boldsymbol{\theta}}\right)\left(1+r_{i}^{-1}\boldsymbol{U}_{i}^{\top}\hat{\boldsymbol{\theta}}-2^{-1}r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2}+\delta_{1i}\right)=0$
where $\delta_{1i}=O_{p}\big{\{}\big{(}r_{i}^{-1}\boldsymbol{U}_{i}^{\top}\hat{\boldsymbol{\theta}}-2^{-1}r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2}\big{)}^{2}\big{\}}=O_{p}\left(n^{-1}\right)$. Then, we have
$n^{-1}\sum_{i=1}^{n}\left(1-2^{-1}r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2}+\delta_{1i}\right)\boldsymbol{U}_{i}+n^{-1}\sum_{i=1}^{n}r_{i}^{-1}\left(\boldsymbol{U}_{i}^{\top}\hat{\boldsymbol{\theta}}\right)\boldsymbol{U}_{i}$
$=$ $n^{-1}\sum_{i=1}^{n}r_{i}^{-1}\left(1-2^{-1}r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2}+\delta_{1i}\right)\hat{\boldsymbol{\theta}}+n^{-1}\sum_{i=1}^{n}r_{i}^{-2}\left(\boldsymbol{U}_{i}^{\top}\hat{\boldsymbol{\theta}}\right)\hat{\boldsymbol{\theta}},$
which implies
$n^{-1}\sum_{i=1}^{n}\left(1-2^{-1}r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2}+\delta_{1i}\right)\boldsymbol{U}_{i}+n^{-1}\sum_{i=1}^{n}r_{i}^{-1}\left(\boldsymbol{U}_{i}^{\top}\hat{\boldsymbol{\theta}}\right)\boldsymbol{U}_{i}=$ $n^{-1}\sum_{i=1}^{n}r_{i}^{-1}\left(1+\delta_{1i}+\delta_{2i}\right)\hat{\boldsymbol{\theta}},$
where $\delta_{2i}=r_{i}^{-1}\boldsymbol{U}_{i}^{\top}\hat{\boldsymbol{\theta}}-2^{-1}r_{i}^{-2}\|\hat{\boldsymbol{\theta}}\|_{2}^{2}=O_{p}(\delta_{1i}^{1/2})$. Then, we obtained that
$\hat{\boldsymbol{\theta}}=\left\{\zeta_{1}+O_{p}\left(\zeta_{1}n^{-1/2}\right)\right\}^{-1}\Big{(}n^{-1}\sum_{i=1}^{n}\mathbf{U}_{i}+\varrho\Big{)}=n^{-1}\zeta_{1}^{-1}\sum_{i=1}^{n}\mathbf{U}_{i}+O\left(\zeta_{1}^{-1}\varrho\right),$
where $\|\varrho\|_{2}=O_{p}(n^{-1})$.
S2.2 Proof of Theorem 2.1
@@
Thus, according to Lemma S2.3, we have
@@
By lemma S2.4, we have, there exist a positive constant $C_{\alpha}$ such that $\left\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\right\|_{2}^{2}\leq C_{\alpha}^{2}\zeta_{1}^{-2}n^{-1}$ with probability larger than $61-\alpha/$ 6. Additionally, by Chebyshev inequality,
@@
By Lemma S2.4, we can write $\hat{\boldsymbol{\theta}}=\zeta_{1}^{-1}\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{U}_{i}+\zeta_{1}^{-1}\varrho$, where $\|\varrho\|_{2}=O_{p}(n^{-1})$. So
@@
So $P\left(\|B_{1}\|{{}_{2}}\geq 3\sqrt{2}\alpha^{-1/2}n^{-1}\sigma_{v}\|{\bf S}\|_{2}\right)\leq\frac{\alpha}{18}.$ Additionally,
@@
S2.4 Proof of Theorem 3.1
LEMMA S2.5
@@
Proof of Theorem 3.1: Note that $U(\boldsymbol{X})$ has the same distribution as $S(X)=\frac{\boldsymbol{X}-\tilde{\boldsymbol{X}}}{\|\boldsymbol{X}-\tilde{\boldsymbol{X}}\|{{}_{2}}}$ where $\boldsymbol{X},\tilde{\boldsymbol{X}}\sim EC_{d}(\boldsymbol{\mu},{\bf\Sigma},\xi)$ and are independent. By Lemma B.4 in Han and Liu (2018), for any $\boldsymbol{v}=\left(v_{1},\ldots,v_{d}\right)^{T}\in\mathbb{S}^{d-1}$, Equation (10) holds with
@@
where $\left(Y_{1},\ldots,Y_{d}\right)^{T}\sim N_{d}\left(\mathbf{0},\mathbf{I}_{d}\right)$ is standard Gaussian. Thus, by Lemma S2.5, when $(s\log(ed/s)+\log(1/\alpha))/n\rightarrow 0$, with probability at least $1-2\alpha$, we have
$\|\widetilde{\mathbf{S}}-\mathbf{S}\|_{2,s}\leq$ $C_{0}\left(\sup_{\boldsymbol{v}\in\mathbb{S}^{d-1}}2\left\|\boldsymbol{v}^{T}U(\boldsymbol{X})\right\|_{\psi_{2}}^{2}+\|\mathbf{S}\|_{2}\right)^{1/2}\sqrt{\frac{s(3+\log(d/s))+\log(1/\alpha)}{n}}.$
Taking the same procedure as Lemma S2.4, we can have $\|\hat{\boldsymbol{\theta}}\|_{2,s}=O_{p}(\sqrt{\frac{s}{d}}\zeta_{1}^{-1}n^{-1/2})$. So taking the same procedure as the proof of Theorem 2.1, we have, there exist some positive constant $C_{1}>0$,
@@
for sufficient large $n$ and with probability larger than $1-\alpha$. Then, by the triangle inequality, we obtain the first result. Specially, if $\operatorname{rank}(\boldsymbol{\Sigma})=q$ and $\left\|\boldsymbol{u}_{1}(\boldsymbol{\Sigma})\right\|_{0}\leq s$, by Theorem 4.2 in Han and Liu (2018), we have